40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9013     Entry time: Thu Aug 15 09:34:12 2013
Author: Steve 
Type: Update 
Category: General 
Subject: Wilcoxon cables rescued 

Eric and Steve,

 We removed Wilcoxon Accelerometer PS and Amplifier unit under the BS optical tabel yesterday. The six cabels going to DAQ  were labeled and left in place. Gain setting were 100, except channel 3 was 10.

The ~ 40 m long 2 sets of 3 cables were very happy to get their kinks out. Especially the set going just south of ITMX optical  table.

We have to take better care of these cables! Your data will be useless this way.

Attachment 1: rescuedGraycables.jpg  225 kB  | Hide | Hide all
rescuedGraycables.jpg
Attachment 2: wilconoxOut.jpg  72 kB  | Hide | Hide all
wilconoxOut.jpg
Attachment 3: chanGains.jpg  64 kB  | Hide | Hide all
chanGains.jpg
ELOG V3.1.3-